OpenStudy (anonymous):

Find the volume of each right prism with the given features. a. The bases are equilateral triangles with sides of length 8 ft, and the height of the prism is 10ft. b.The bases are trapezoids with bases of lengths 7 cm and 9 cm perpendicular to one side of length 6 cm, and the height of the prism is 12cm.

OpenStudy (anonymous):

The answer actually in book equaled this; a=277 b=576

OpenStudy (anonymous):

Sorry I calculated the surface area instead of volume..... Wait I'll calculate the volume......

OpenStudy (anonymous):

a) The bases are equilateral triangles with sides of length 8 ft, and the height of the prism is 10ft. Volume = area of base x height base is equilateral triangle so sqrt(3) 1.73 1.73 area of base = ------- * (side)^2 = ----- * (8 ft)^2 = ----- * 64 ft^2 = 27.68 ft^2 4 4 4 so Volume = 27.68 ft^2 * 10 ft = 276.8 ft^3 ≈ 277 ft^3 so your answer is a) 277 ft^3